FAR-INFRARED SYNCHROTRON-BASED SPECTROSCOPY OF FURAN: ANALYSIS OF THE ν_{14} − ν_{11} PERTURBATION AND THE ν_{18} AND ν_{19} LEVELS

\underline{D. W. TOKARYK}, S. D. CULLIGAN^{a}, Department of Physics and Centre for Laser, Atomic and Molecular Sciences, University of New Brunswick, Fredericton, NB, Canada E3B 5A3; B. E. BILLINGHURST, Canadian Light Source, Inc., 101 Perimeter Road, University of Saskatchewan, Saskatoon, SK, Canada S7N 0X4; and J. A. van WIJNGAARDEN, Department of Chemistry, University of Manitoba, Winnipeg, MB, Canada R3T 2N2.

The ν_{14} vibrational level of furan lies 603 cm^{−1} above the ground vibrational state. It is the lowest lying vibrational level for which a transition from the ground state is allowed. Other groups have conducted rotational analyses on fundamental bands of furan at 745 cm^{−1} (ν_{13})^{b}, 995 cm^{−1} (ν_{7})^{c} and at 1067 cm^{−1} (ν_{6})^{d}. We have taken the rotationally resolved spectrum of the c-type ν_{14} band at the Canadian Light Source synchrotron with a Bruker IFS125HR Fourier transform spectrometer operating at 0.00096 cm^{−1} resolution, and have found it to be perturbed by the ν_{11} band at 600 cm^{−1}, for which transitions from the ground vibrational state are forbidden. By taking the spectra of the b-type ν_{18} fundamental band and of the very weak c-type ν_{18}- ν_{11} band we have been able to analyze the ν_{14}- ν_{11} perturbation. We have also analyzed the spectrum of the b-type ν_{19} fundamental band.

aCurrent address: Inorganic Chemistry Laboratory, South Parks Road, University of Oxford, UK OX1 3QR

bB. Pankoke, K. M. T. Yamada, G. Winnewisser, Z. Naturforsch. A 48 (1993) 1193–1202.

cA. Mellouki, M. Herman, J. Demaison, B. Lemoine, L. Margulès, J. Mol. Spectrosc. 198 (1999) 348–357.

dA. Mellouki, J. Vander Auwera, J. Demaison, M. Herman, J. Mol. Spectrosc. 209 (2001) 136–138.